A small city with 3 neighborhoods is trying to determine how many firefighters to hire. The marginal benefit that each neighborhood receives in dollars per hour from hiring additional firefighters is given in the table below:

Number of Firefighters

MB of Area A

MB of Area B

MB of Area C

1

$18

$15

$8

2

$14

$12

$6

3

$10

$9

$4

4

$6

$6

$2

5

$2

$3

$0

If a firefighter’s wage is set at $30 per hour, what is the optimal number of firefighters for the city to hire?

2

5

3

1

4


PLEASE BOLD THE CORRECT ANSWER!
Answer #1
Number of firefighters MB of Area A MB of area B MB of area C TB of Area A TB of Area B TB of Area C Total cost Total benefit of all three area Benefit-cost
1 18 15 8 18 15 8 =A2*30 =SUM(E2:G2) =I2-H2
2 14 12 6 =E2+B3 =F2+C3 =G2+D3 =A3*30 =SUM(E3:G3) =I3-H3
3 10 9 4 =E3+B4 =F3+C4 =G3+D4 =A4*30 =SUM(E4:G4) =I4-H4
4 6 6 2 =E4+B5 =F4+C5 =G4+D5 =A5*30 =SUM(E5:G5) =I5-H5
5 2 3 0 =E5+B6 =F5+C6 =G5+D6 =A6*30 =SUM(E6:G6) =I6-H6
Number of firefighters MB of Area A MB of area B MB of area C TB of Area A TB of Area B TB of Area C Total cost Total benefit of all three area Benefit-cost
1 18 15 8 18 15 8 30 41 11
2 14 12 6 32 27 14 60 73 13
3 10 9 4 42 36 18 90 96 6
4 6 6 2 48 42 20 120 110 -10
5 2 3 0 50 45 20 150 115 -35

Answer: 2